OpenStudy (anonymous):

find the exact value of cos(15) and cos(7,5)

OpenStudy (anonymous):

\[\text{Cos}[15{}^{\circ}]=\frac{1+\sqrt{3}}{2 \sqrt{2}} \]Don't believe Cos(7.5) has an exact value.